# Dynamical and structural properties of an absorbing phase transition: a case study from granular systems

### v1

@@ Appendix E Expression for the transport coefficients and numerical values
As mentioned in the main text, we use the simple Enskog’s equilibrium transport coefficients. They should give reasonable results in the active state. Moreover, while our simulations are mostly done at $\phi=0.1$, the theory works well at higher densities (see Fig. 14, where $\phi=0.5$ for instance). Here, we provide a list of the expressions used. They are derived assuming molecular chaos and a Gaussian velocity distribution. We will also use the equilibrium contact value of the pair correlation function [127]:
@@
The pressure is given by [134]:
@@
The transport coefficients we used are the equilibrium ones from Enskog’s theory [179]:
@@
For the transport coefficients taking into account $\Delta$ and $\alpha$, see Ref. 134. The transport coefficient with $\Delta+\gamma$ are not known and should be derived from the Chapman-Enskog procedure.
@@ Appendix F Adiabatic slaving of the temperature field: $\bm{M}^{(r)}$ and theoretical correlation functions
As already stated in the main text, we can find exact theoretical expressions for the eigenvalues of $\bm{M}$ or the dynamic and static correlation functions through Eqs. (47) and (46). However, their expressions are not particularly useful since they are third order polynomial roots or made of ratio of polynomials of degree up to $k^{6}$. Therefore, it becomes necessary to perform approximations to make the computations more tractable and the results clearer. One way is to integrate the temperature field by assuming that it varies quicker than any other field. This assumption is exact at $k=0$ since the others field are conserved and hence, have a relaxation timescale diverging with $1/k$. However, at larger $k$, this assumption will not be as good. Moreover, when $\gamma\neq 0$, the momentum field is also a fast field at small $k$. In what follow, we will only integrate the temperature field following Ref. 92, in order to obtain a hydrodynamic description with only the velocity and density fields.
@@
with $\Gamma=-G_{T}+\kappa_{0}k^{2}/n_{0}>0$ is an inverse relaxation time for the temperature field and where we assumed $\delta T(-\infty)=0$ without loss of generality since we are only concerned about the long time limit. We could then replace this solution for $\delta T(t)$ in the hydrodynamic equation given by $\dot{\bm{\Psi}}=\bm{M}\bm{\Psi}$. We would have eliminated the temperature field exactly, at the price of having to deal with a non-Markovian system with an exponential kernel. At that point, it is possible to indeed follow this route, integrate by parts the equations and eventually perform an expansion on the resulting equation for the density and velocity. However, we follow Ref. 92 and immediately perform the adiabatic expansion on Eq. (64) by performing successive integration by parts to expand the integral in powers of $\Gamma^{-1}$. $\Gamma$ is assumed to be large compared to the other relaxation scale thanks to its finiteness at $k=0$, allowing us to truncate the expansion without significant loss of physical accuracy. The expansion is performed as follows:
[…]
@@
This reduced hydrodynamics model includes correction to the simple model introduced by Lei and Ni in Ref. 93 to describe a hyperuniform fluid. The first stability criterion is again apparent here in the velocity-density coupling. Similar systems, with an effective $\Delta$, were found to be unstable due to the positivity of this term [213, 214]. At equilibrium, spinodal instabilities related to phase transitions are also captured by the term $\bm{M}_{u^{\parallel}\rho}$ in a hydrodynamic theory, since the compressibility $1/(\rho p_{\rho})$ becomes negative. However, such instability analysis on the hydrodynamic matrix cannot capture the binodals, metastability and coexistence densities, as these require thermodynamic information or additional constraints [215, 216, 217, 218, 219].
We now turn to correlation computation. Due to the peculiar correlation of the derivative of the white noise, it is not very practical to use the relations used in the main text to find $\bm{S}(\bm{k},w)$ and $\bm{S}(\bm{k})$ (Eqs. (46) and (47)). Moreover, the adiabatic expansion in all its glory starts to become heavy to carry since second derivative of noises arise. In order to easily compute $\bm{S}(\bm{k},w)$ and $\bm{S}(\bm{k})$, we chose to ignore the time derivative contribution $\dot{S}(t)$ in the expansion of $\delta T(t)$ which amounts to simply setting $\delta\dot{T}=0$. This is inconsistent but leads to good results. For simplicity and for illustrative purposes, we will work in the case $\gamma=0$ which implies $G_{\rho}=0$. This allows us to nicely predict the observed shape of the structure factor for the system “$\Delta$ alone” in Fig. 10.
Figure 14: Comparison between the structure factor found in the simulations of the simple $\Delta$ model without damping, the full theory (obtained from the full hydrodynamic matrix) and the adiabatic slaving (Eq. (74)). $\alpha=0.9$, $\Delta/(\sigma\hat{\tau})=0.05$, $\phi=0.5$ and $N=2.5\times 10^{6}$.
@@
We can integrate between $\infty$- ∞ and $\infty$ and divide by $2\pi$ Eq. (71) to find the static correlation function. This gives the simple result:
@@
Interestingly, the equilibrium limit $G_{T}\to 0$, which occurs, for example, when $\alpha\to 1$, is singular. Indeed, as we approach this singular equilibrium limit $\alpha\to 1$, the value of $S(0)$ decreases to 00, moving us further away from the true equilibrium result proportional to the temperature and the compressibility. However, as this limit is approached, increasingly smaller values of $k$ are required for the transient hyperuniform scaling to manifest. This implies that infinitely close to the equilibrium limit, an infinitely large system is required to observe the non-equilibrium behavior of the structure factor. One might argue that our theoretical predictions become unreliable near equilibrium because the parameter $T$ does not vary rapidly. However, it is important to note that there will always exist a value of $k$ for which the conserved fields $\rho$ and $\bm{u}$ evolve more slowly than the non-conserved temperature field. Therefore, the asymptotic behavior remains reliable and, indeed, exact.

### v2

@@ Appendix E Expression for the transport coefficients and numerical values
As mentioned in the main text, we use the simple Enskog’s equilibrium transport coefficients. They should give reasonable results in the active state. Moreover, while our simulations are mostly done at $\phi=0.1$, the theory works well at higher densities (see Fig. 14, where $\phi=0.5$ for instance). Here, we provide a list of the expressions used. They are derived assuming molecular chaos and a Gaussian velocity distribution. We will also use the equilibrium contact value of the pair correlation function [mulero2009equation]:
@@
The pressure is given by [garzo2018enskog]:
@@
The transport coefficients we used are the equilibrium ones from Enskog’s theory [gradenigo2011fluctuating]:
@@
For the transport coefficients taking into account $\Delta$ and $\alpha$, see Ref. garzo2018enskog. The transport coefficient with $\Delta+\gamma$ are not known and should be derived from the Chapman-Enskog procedure.
@@ Appendix F Adiabatic slaving of the temperature field: $\bm{M}^{(r)}$ and theoretical correlation functions
As already stated in the main text, we can find exact theoretical expressions for the eigenvalues of $\bm{M}$ or the dynamic and static correlation functions through Eqs. (47) and (46). However, their expressions are not particularly useful since they are third order polynomial roots or made of ratio of polynomials of degree up to $k^{6}$. Therefore, it becomes necessary to perform approximations to make the computations more tractable and the results clearer. One way is to integrate the temperature field by assuming that it varies quicker than any other field. This assumption is exact at $k=0$ since the others field are conserved and hence, have a relaxation timescale diverging with $1/k$. However, at larger $k$, this assumption will not be as good. Moreover, when $\gamma\neq 0$, the momentum field is also a fast field at small $k$. In what follow, we will only integrate the temperature field following Ref. brito2013hydrodynamic, in order to obtain a hydrodynamic description with only the velocity and density fields.
@@
with $\Gamma=-G_{T}+\kappa_{0}k^{2}/n_{0}>0$ is an inverse relaxation time for the temperature field and where we assumed $\delta T(-\infty)=0$ without loss of generality since we are only concerned about the long time limit. We could then replace this solution for $\delta T(t)$ in the hydrodynamic equation given by $\dot{\bm{\Psi}}=\bm{M}\bm{\Psi}$. We would have eliminated the temperature field exactly, at the price of having to deal with a non-Markovian system with an exponential kernel. At that point, it is possible to indeed follow this route, integrate by parts the equations and eventually perform an expansion on the resulting equation for the density and velocity. However, we follow Ref. brito2013hydrodynamic and immediately perform the adiabatic expansion on Eq. (64) by performing successive integration by parts to expand the integral in powers of $\Gamma^{-1}$. $\Gamma$ is assumed to be large compared to the other relaxation scale thanks to its finiteness at $k=0$, allowing us to truncate the expansion without significant loss of physical accuracy. The expansion is performed as follows:
$\begin{split}\delta T(t)&=\int_{-\infty}^{t}\text{d}t^{\prime}e^{\Gamma(t^{\prime}-t)}\mathcal{S}(t^{\prime})\\ &=\left[\dfrac{e^{\Gamma(t^{\prime}-t)}\mathcal{S}(t^{\prime})}{\Gamma}\right]_{-\infty}^{t}-\dfrac{1}{\Gamma}\int_{-\infty}^{t}\text{d}t^{\prime}e^{\Gamma(t^{\prime}-t)}\mathcal{S}^{\prime}(t^{\prime})\\ &=\sum_{n=0}^{N}\dfrac{(-1)^{n}}{\Gamma^{n+1}}\dfrac{\partial^{n}\mathcal{S}(t)}{\partial t^{n}}+\\ &~\sum_{m=N+1}^{\infty}\dfrac{(-1)^{m}}{\Gamma^{m}}\int_{-\infty}^{t}dt^{\prime}e^{\Gamma(t^{\prime}-t)}\dfrac{\partial^{m}\mathcal{S}(t)}{\partial t^{m}}\\ &=\dfrac{\mathcal{S}(t)}{\Gamma}-\dfrac{\dot{\mathcal{S}}(t)}{\Gamma^{2}}+\mathcal{O}(\Gamma^{-3})\end{split}$ (65)
@@
This reduced hydrodynamics model includes correction to the simple model introduced by Lei and Ni in Ref. lei2019hydrodynamics to describe a hyperuniform fluid. The first stability criterion is again apparent here in the velocity-density coupling. Similar systems, with an effective $\Delta$, were found to be unstable due to the positivity of this term [risso2018effective, argentina2002van]. At equilibrium, spinodal instabilities related to phase transitions are also captured by the term $\bm{M}_{u^{\parallel}\rho}$ in a hydrodynamic theory, since the compressibility $1/(\rho p_{\rho})$ becomes negative. However, such instability analysis on the hydrodynamic matrix cannot capture the binodals, metastability and coexistence densities, as these require thermodynamic information or additional constraints [chiu2024theorynonequilibriummulticomponentcoexistence, solon2015pressure, solon2018generalized, omar2023mechanical, feng2024theoryanomalousphasebehavior].
We now turn to correlation computation. Due to the peculiar correlation of the derivative of the white noise, it is not very practical to use the relations used in the main text to find $\bm{S}(\bm{k},w)$ and $\bm{S}(\bm{k})$ (Eqs. (46) and (47)). Moreover, the adiabatic expansion in all its glory starts to become heavy to carry since second derivative of noises arise. In order to easily compute $\bm{S}(\bm{k},w)$ and $\bm{S}(\bm{k})$, we chose to ignore the time derivative contribution $\dot{S}(t)$ in the expansion of $\delta T(t)$ which amounts to simply setting $\delta\dot{T}=0$. This is inconsistent but leads to good results. For simplicity and for illustrative purposes, we will work in the case $\gamma=0$ which implies $G_{\rho}=0$. This allows us to nicely predict the observed shape of the structure factor for the system $\Delta$ alone” in Fig. 10.
Figure 14: Comparison between the structure factor found in the simulations of the simple $\Delta$ model without damping, the full theory (obtained from the full hydrodynamic matrix) and the adiabatic slaving (Eq. (74)). $\alpha=0.9$, $\phi=0.5$ and $N=2.5\times 10^{6}$.
@@
We can integrate between -$\infty$ and $\infty$ and divide by $2\pi$ Eq. (71) to find the static correlation function. This gives the simple result:
@@
Interestingly, the equilibrium limit $G_{T}\to 0$, which occurs, for example, when $\alpha\to 1$, is singular. Indeed, as we approach this singular equilibrium limit $\alpha\to 1$, the value of $S(0)$ decreases to 0, moving us further away from the true equilibrium result proportional to the temperature and the compressibility. However, as this limit is approached, increasingly smaller values of $k$ are required for the transient hyperuniform scaling to manifest. This implies that infinitely close to the equilibrium limit, an infinitely large system is required to observe the non-equilibrium behavior of the structure factor. One might argue that our theoretical predictions become unreliable near equilibrium because the parameter $T$ does not vary rapidly. However, it is important to note that there will always exist a value of $k$ for which the conserved fields $\rho$ and $\bm{u}$ evolve more slowly than the non-conserved temperature field. Therefore, the asymptotic behavior remains reliable and, indeed, exact.